OpenStudy (anonymous):

write each polynomial in factored form 2x3+10x2+12x

OpenStudy (anonymous):

Help The exponent is after the Xs

OpenStudy (triciaal):

first factor out an x then do the quadratic

OpenStudy (anonymous):

how

OpenStudy (anonymous):

x2

OpenStudy (anonymous):

subtract x2 in each one

OpenStudy (triciaal):

when you divide each term by x what do you get?

OpenStudy (anonymous):

2x2+10x2+12

OpenStudy (triciaal):

you can also divide by 2

OpenStudy (anonymous):

oooooo

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

5x+6

OpenStudy (triciaal):

so you have 2x(x^2 + 5x + 6)

OpenStudy (anonymous):

still a little confused

OpenStudy (anonymous):

as to how u would show it

OpenStudy (triciaal):

2x(x +3)(x+2)

OpenStudy (anonymous):

thanks I have quiz tomorrow

OpenStudy (triciaal):

divide each term you just forgot

OpenStudy (triciaal):

think about long division

OpenStudy (triciaal):

2x^3 /2x = x^2

OpenStudy (anonymous):

Find the zeros of the function. State the multiplicty of multiple zeros y=(x+4)^2(x+1)

OpenStudy (anonymous):

y=(x^2+16)(x+1)

OpenStudy (triciaal):

(a + b)^2 =a^2 + 2ab+ b^2 (x+4)^2 =